OpenStudy (anonymous):

Write an equation of a line with slope m = -2/ 5 and passing through the point (5, -2). A. y = -2/5 x - 2 B. y = -2/5 x C. y = -2/5 x + 2 D. y = -2/5 x + 4

OpenStudy (kkutie7):

so you want to use this equation \[y=m(x-x)+y\]

OpenStudy (kkutie7):

or better \[y=m(x-x_{1})+y_{1}\]

OpenStudy (kkutie7):

you have m and \[x_{1}=5\] \[y-{1}=-2\]

OpenStudy (kkutie7):

*\[y_{1}=-2\]

OpenStudy (anonymous):

y=mx+b

OpenStudy (kkutie7):

\[y=\frac{-2}{5}(x-5)-2\] can you do it from here?

OpenStudy (anonymous):

so y=-2/5x but idk the rest. im kinda confused @Kkutie7

OpenStudy (kkutie7):

Ok so use the equation I gave you and multiply through here look \[y=\frac{-2}{5}x+\frac{10}{5}-2\] \[y=\frac{-2}{5}x+2-2\] so its just basically the slope.

OpenStudy (anonymous):

so y= -2/5x+2-2 is the answer?

OpenStudy (kkutie7):

yes and no =) you can 2-2=0 so its just y=-2/5x